### v1

# Guangyu Zhu, Yiqin Deng, , Xianhao Chen, , Haixia Zhang, , Yuguang Fang, , Tan F. Wong

Abstract
Federated learning (FL) allows multiple parties (distributed devices) to train a machine learning model without sharing raw data. How to effectively and efficiently utilize the resources on devices and the central server is a highly interesting yet challenging problem. In this paper, we propose an efficient split federated learning algorithm (ESFL) to take full advantage of the powerful computing capabilities at a central server under a split federated learning framework with heterogeneous end devices (EDs). By splitting the model into different submodels between the server and EDs, our approach jointly optimizes user-side workload and server-side computing resource allocation by considering users’ heterogeneity. We formulate the whole optimization problem as a mixed-integer non-linear program, which is an NP-hard problem, and develop an iterative approach to obtain an approximate solution efficiently. Extensive simulations have been conducted to validate the significantly increased efficiency of our ESFL approach compared with standard federated learning, split learning, and splitfed learning.
Index Terms:
Distributed machine learning, federated learning, split learning, wireless networking.
I Introduction
In recent years, machine learning (ML) has attracted intensive attention in many fields and numerous artificial intelligence (AI) applications, such as computer vision, smart health, connected and autonomous driving, information access control, and security surveillance [1]. According to Cisco [2], there were nearly $850$ zettabyte of data generated by people, machines, and things at the network edge in $2021$. It is definitely infeasible to simply send this huge data volume to a central server to process or compute. To do so, a tremendous network bandwidth is required, incurring intolerable latency. Hence, distributed machine learning algorithms have been developed to cope with the aforementioned challenges for large geographically distributed volume of data by distributing the ML workloads to EDs [3]. Moreover, awareness and concerns about users’ privacy have been raised in the digitalized world [4]. Following the Privacy-by-Design (PbD) principle, the best way to achieve user privacy is not to disclose raw data, and so it would be more effective for EDs not to share private raw data as much as possible during the machine learning process.
Federated learning (FL), a privacy-preserving distributed ML technique enables EDs to collaboratively learn a global ML model without sharing their raw data, consequently reducing the requirement of communication bandwidth between EDs and a central server as well. The intuitive idea of privacy-preserving distributed ML was investigated independently by some researchers Xu et al. [5, 6, 7], Shokri et al. [8], Gong et al. [9]. McMahan et al. [10] first constructed a distributed ML framework based on decentralized datasets held by different users, and claimed their algorithm as FL. The original FL algorithm updates ML models on EDs locally, and aggregates the updated ML models to obtain the global model at the central server remotely, in contrast to the traditional ML requiring EDs to upload their private data to where the ML model is trained. Since FL retains users’ private data on the EDs without sharing raw data with the server, the server only aggregates users’ locally updated models, as the privacy of end users is naturally preserved to some extent. Besides, as all training processes are performed by the EDs, the computation and communication capacities of EDs can also be utilized. FL provides various advantages, such as data privacy preservation, reduced communication latency, and enhanced learning performance [2]. However, pushing all training workload to EDs sometimes is impracticable. Training complex ML models often consumes unacceptable amount of training memory and computing power on Internet of Things (IoT) devices with limited communication and computing resources, and incurs intolerable latency. Besides, the ML model size sometimes reaches GBs (e.g., GPT-1 has 0.12 billion parameters, GPT-2 has 1.5 billion parameters, whereas GPT-3 has more than 175 billion parameters), resulting in a significant communications burden duo to the need of frequently uploading local ML models.
To cope with the dilemma between insufficient ED resources and complicated ML models, we leverage another ML technique, called split federated learning (SFL) [11], which introduces model splitting from split learning (SL) [12] to FL. The split federated learning framework was proposed by Thapa et al. [11] in 2020 to integrate federated learning with split learning, shown in Figure 1. Under SFL framework, the main server helps the training process for each client, and the Fed server is tasked with the aggregation of all locally updated models. The SFL approach presents a compelling advantage for resource-constrained environments, since the main server undertake parts of local training workload.
Figure 1: The architecture of splitfed learning (SFL) system.
From the conventional FL and SFL implementations across heterogeneous devices, we have identified several limitations that can significantly impede the system’s overall performance and efficiency. In synchronous FL and SFL, the aggregation process is inherently constrained by the pace of the slowest participant due to the necessity for the server to collect updated local gradients from all selected EDs. This synchronicity results in a scenario where devices with abundant computational resources are rendered idle as they wait for the transmission of local models from less capable devices, often referred to as “straggler”. This inefficiency is primarily attributed to the varying and unpredictable communication and computational capabilities of the heterogeneous EDs. Therefore, addressing resource heterogeneity (RH), also called the system heterogeneity problem in FL, is critical to enhancing the efficiency and overall feasibility of SFL.
In this paper, we propose a novel efficient split federated learning algorithm (ESFL) to boost training efficiency and performance considering the privacy-preserving constraint. The cornerstone of ESFL lies in its strategic utilization of heterogeneity in system resources and device capabilities. Unlike standard synchronous FL and SFL, our approach involves the server in the training phase but dynamically adjusts the distribution of user-side workload and server-side resources. The ESFL algorithm thereby capitalizes on the intrinsic resource variability across EDs to optimize ML outcomes and system efficiency.
Our major contributions in this paper are summarized as follows.
  • 1)
    We design ESFL, a novel distributed machine learning framework, which significantly improves the training efficiency of SFL by taking ED heterogeneity into consideration.
  • 2)
    We formulate a mixed-integer non-linear program (MINLP) by jointly considering the allocation of user-side workload (model separation) and server-side resource, and develop an iterative optimization algorithm to find a suboptimal solution with a low time complexity.
  • 3)
    We evaluate the performance of our ESFL approach through extensive simulations compared to the state-of-the-art methods such as FL, SL and SFL, and demonstrate the superiority of the proposed ESFL framework.
The remainder of this paper is organized as follows. In Section II, we discuss related works about FL, SFL, and RH. In Section III, we expound on the framework and system model of ESFL. In Section IV, we present the optimization problem formulation and the solution to the joint resource allocation and model separation problem. In Section V, results of extensive simulations and experiments are presented. In Section VI, we summarize the proposed algorithm and experimental results and conclude the paper.
II Related Works
There exist quite extensive research on splitting learning and federated learning in the current literature. In this section, we will concentrate on closely related works to review. In [13], Ang et al. offered a robust architecture for FL to increase communication efficiency by reducing transmission noise in wireless networks. To address the communications overhead in FL, Wang et al. [14] developed an algorithm, named Communication Mitigated Federated Learning (CMFL), to eliminate irrelevant local model updates that were trained over users’ biased data. It should be noted that the aforementioned two methods only address communication efficiency in FL, as all training processes are executed by the user-side EDs. FedMMD was proposed in [15] to reduce communication and computing costs during the local training process in FL using a two-stream model with Maximum Mean Discrepancy (MMD) to replace the local training for a single model in FedAVG [10]. However, while this method reduces the number of communication rounds, it concurrently increases the user-side workload due to the need of computing the MMD loss. Wenqi et al. [16] address both communication and computing resource heterogeneities in wireless FL by providing a joint device scheduling and resource allocation strategy. Despite this integrated approach, the limitation of FL is still evident, as the user-side workload cannot be reduced.
In [11], Thapa et al. designed a novel framework, called splitfed learning (SFL), to take advantage of parallel training among different users in FL and the model splitting in SL to reduce the computing workload on EDs. In [17], Gao et al. implemented SFL and evaluated the performance on IoT devices. In [18], Lin et al. developed an efficient parallel split learning algorithm by applying the last-layer gradient aggregation to reduce communication and computing overheads in SL. In [19], Wu et al. designed a resource allocation strategy for cluster-based SFL. In [20], Kim et al. devised a bargaining game to negotiate the cut layer in personalized parallel SL. All aforementioned literature on SFL overlook the crucial aspect of the joint consideration of both user-side resource and workload heterogeneities. This oversight is evidenced by the fact that the cut layer remains the same at all EDs. However, it is obvious that adjusting cut layers can significantly change user-side workload distribution.
To reduce the computing and communication workload at EDs when considering RH in FL, Sattler et al.  [21] designed a novel model compression algorithm to extend the commonly used top-k gradient sparsification to FL to compress both model uploading and downloading. However, since model compression inevitably results in performance degradation, it is advisable to maintain the integrity of the model architecture throughout the training process. With the emergence of edge computing [22], utilizing the computing resources at the both the central server and edges can be leveraged to reduce workload at EDs. For instance, in [23], Wang et al. proposed to enable EDs to collaborate with edge nodes by exchanging model parameters to reduce the user-side workload, and to apply deep reinforcement learning to optimize the operations of mobile edge computing, caching, and communications. Several researchers attempted to leverage edge nodes to assist FL by fully uploading local training tasks to trusted edge nodes to reduce the user-side workload [24]. Unfortunately, the trustworthiness of training edge nodes is often hard to guarantee. Our ESFL algorithm introduce an integrated strategy for the allocation of server-side computing resource and user-side training workload. This approach is designed to accommodate the variations in computing and communication resources inherent in heterogeneous EDs. We will present the detailed design next.
III Efficient Split Federated Learning
III-A Motivation
While SFL only address the resource constraints at EDs, it does not account for the variations in data distribution (DH) and resource availability (resource heterogeneity or RH) inherent in EDs. Within a SFL framework, clients or EDs are required to partition their models following an identical structure, leading to a scenario where the communication and computational workload on the client side is influenced solely by the volume of data. In contrast, our Efficient Split Federated Learning (ESFL) takes a holistic approach, considering both the client-side workload and the server-side computing resource allocation, which is designed to mitigate “stragglers” problem in FL, thereby enhancing the training efficiency of SFL.
III-B ESFL framework
In this subsection, we elaborate our framework of ESFL consist of four components, namely, split training, federated aggregation, communication model, and resource allocation. We assume that an ML task is composed of multiple training rounds. The FL server first initializes ML model in the initial round of training and the subsequent training rounds. The one-round training procedure is given as follows.
  • 1)
    User selection: The server selects users randomly from the available users (EDs), who are willing and ready to participate this round of training (these users are called the selected users).
  • 2)
    Model splitting and resource allocation: The server first acquires the information of the selected EDs, including data amount, and available communication and computing resources. Then, the server jointly splits the ML model into user-side models and server-side models (the so-called cut layer decision), and allocates adequate server-side resources based on the users’ information and the ML model structure for the selected user.
  • 3)
    Model distribution: The server distributes user-side models (with the corresponding architectures) to the selected users.
  • 4)
    Split training: The server and all selected users then collaboratively update both user-side and server-side ML models simultaneously and the server determines the ML hyperparameters, such as learning rate, data batch size, and local training epochs.
  • 5)
    Federated aggregation: After repeating several epochs of the split training step, all selected users that have finished their training upload their updated user-side ML models to the server, and then the server generates an updated global ML model based on the user-side ML models collected from the selected users and the corresponding server-side ML models at the server.
¡¡¡¡¡¡¡ HEAD
III-B1 Split Training
Different from federated learning, which only relies on EDs to update local ML models, we split the local training across EDs and the server. Similar to the SplitFed [11], local training is repeatedly conducted $\epsilon_{i}$ epochs for user $i$ before one-round global ML model aggregation at the server. The user-side ML model for user $i$ is $\mathcal{\textbf{W}}_{u,i}^{r,e}$ at epoch $e$ in the $r$-th round. We denote the updated local user-side ML model at the epoch $e+1$ as
$\mathcal{\textbf{W}}_{u,i}^{r,e+1}$ $=BP(\mathcal{\textbf{W}}_{u,i}^{r,e},\rho_{r},\nabla{\textbf{A}}_{l_{i}}^{r,e}),$ (1)
where BP is the backpropagation algorithm, $\rho_{r}$ is the local learning rate in the $r$-th round, and $\nabla{\textbf{A}}_{l_{i}}^{r,e}={\textbf{A}}_{l_{i}}^{r,e}-\hat{\textbf{A}}_{l_{i}}^{r,e}$ is the activation difference, which is the loss value for BP, and $\hat{\textbf{A}}_{l_{i}}^{r,e}$ is the updated activation calculated by the server using backpropagation algorithm with the shared labels and the estimated labels.
The server-side ML model for user $i$ at epoch $e+1$ in the $r$-th round is given by
$\mathcal{\textbf{W}}_{s,i}^{r,e+1}$ $=BP(\mathcal{\textbf{W}}_{s,i}^{r,e},\rho_{r},\nabla\mathcal{l}(\mathcal{\textbf{W}}_{s,i}^{r,e},\hat{\textbf{Y}},\textbf{Y})).$ (2)
To update the server-side ML model, we use the loss function $\mathcal{l}(\mathcal{\textbf{W}}_{s,i}^{r,e},\hat{\textbf{Y}},\textbf{Y})$, where $\hat{\textbf{Y}}$ is the estimated result computed by the server-side ML model and Y is the true label shared by the user. The estimated result $\hat{\textbf{Y}}$ is the output of the current server-side ML model $\mathcal{\textbf{W}}_{s,i}^{r,e}$ and the input data ${\textbf{A}}_{l_{i}}^{r,e}$ is the activation generated by the user $i$’s samples and user-side ML model. Split training divides the local training into four parts in FL, i.e., user-side forward propagation, server-side forward propagation, server-side backpropagation, and user-side backpropagation. The entire split training procedure is shown in Figure 2.
Figure 2: The split training procedure, where all selected users simultaneously transmit activation data and labels to the server, and the server sends back the corresponding activation difference.
The pseudocode for our ESFL is shown in Algorithm 1 and Algorithm 2. The split training is shown in Algorithm 1, while Step $2$ is to let users send activations of the cut layer together with the label to the server. ======= The pseudocode for ESFL is shown in Algorithm 1 and Algorithm 2. The Split training is shown in Algorithm 1, where Step $2$ is to let users send activations of the cut layer together with the label to the server. The detail implementing steps of ESFL are shown in Algorithm 2. ¿¿¿¿¿¿¿ 8fed39b787cde27c064d35637df79221d613d121
Input: At epoch $e$ of round $r$, user-side model $\mathcal{\textbf{W}}_{u,i}^{r,e}$, server-side model $\mathcal{\textbf{W}}_{s,i}^{r,e}$, and local learning rate $\rho$;
Output : Updated user-side $\mathcal{\textbf{W}}_{u,i}^{r+1,e}$ and server-side model $\mathcal{\textbf{W}}_{s,i}^{r+1,e}$;
1 User $i$ forwards propagation of $\mathcal{\textbf{W}}_{u,i}^{r,e}$ and generate activation $\textbf{A}_{l_{i}}^{r,e}$ and the label vector Y;
2 User $i$ sends $\textbf{A}_{l_{i}}^{r,e}$ at the cut layer and Y to the server;
3 The server uses backpropagation to calculate the server-side gradient $\nabla\mathcal{l}(\mathcal{\textbf{W}}_{s,i}^{r,e},\textbf{Y})$ and the activation difference $\nabla{\textbf{A}}_{l_{i}}^{r,e}$;
4 The server-side model update: $\mathcal{\textbf{W}}_{s,i}^{r,e+1}=BP(\mathcal{\textbf{W}}_{s,i}^{r,e},\rho,\nabla\mathcal{l}(\mathcal{\textbf{W}}_{s,i}^{r,e},\textbf{Y}))$;
5 The server sends $\nabla{\textbf{A}}_{l_{i}}^{r,e}$ to user $i$;
The user-side model update: $\mathcal{\textbf{W}}_{u,i}^{r,e+1}=BP(\mathcal{\textbf{W}}_{u,i}^{r,e},\rho,\nabla{\textbf{A}}_{l_{i}}^{r,e})$
Algorithm 1 SplitUpdate
III-B2 Federated Aggregation
¡¡¡¡¡¡¡ HEAD In ESFL, we leverage FedAVG algorithm [10] to aggregate multiple user-side and server-side updated ML models and add a scaling factor $\eta$. We call this aggregation method as federated aggregation, which mitigates the training oscillation by leveraging the global ML model in the previous round, compared with FedAVG, which only aggregates current-round updated local ML model to generate next-round global ML models. The architecture of federated aggregation is shown in Figure 3. We define the whole ML model trained by user $i$ in round $r$ as $\mathcal{w}_{i}^{r}$ and the global model in round $r$ is $\mathcal{\textbf{W}}^{r}$. The update of the global ML model in round ${r+1}$ is ======= In ESFL, we leverage the FedAVG algorithm [10] to aggregate multiple user-side and server-side updated ML models using $\eta$ as a step size factor. We call this aggregation method federated aggregation, which mitigates training oscillations by leveraging the global ML model in the previous round. This differ from FedAVG which only aggregates the current-round updated local ML models to generate the next-round global ML model. The architecture of federated aggregation is shown in Figure 3. We define the whole ML model trained by user $i$ in round $r$ as $\mathcal{w}_{i}^{r}$ and the global model in round $r$ is $\mathcal{\textbf{W}}^{r}$. The update of the global ML model in round ${r+1}$ is ¿¿¿¿¿¿¿ 8fed39b787cde27c064d35637df79221d613d121
$\begin{gathered}\mathcal{\textbf{W}}^{r+1}=\mathcal{\textbf{W}}^{r}-\eta*\left(\mathcal{\textbf{W}}^{r}-\mathcal{\textbf{W}}^{r}_{*}\right)\\ =\mathcal{\textbf{W}}^{r}-\eta*\left(\mathcal{\textbf{W}}^{r}-\sum_{i}\frac{n_{i}*\mathcal{\textbf{W}}_{i}^{r}}{N}\right),\end{gathered}$ (3)
where $N=\sum_{i}n_{i}$ indicating the number of total training samples.
The whole training model is split into two parts, namely, user-side model $\textbf{W}_{u,i}^{r}$ trained by a user-side ED and the server-side model $\textbf{W}_{s,i}^{r}$ trained by a virtual server $v_{i}$. All virtual servers are either virtual machines or containers at the central server, which are allocated with corresponding computing and communication resources according to the model training demands at this round. Thus, after concatenating the user-side ML model and the server-side ML model, for each selected user, the concatenated ML model will have the same architecture as the global ML model. The updated concatenated ML model for user $i$ in round $r+1$ is
$\mathcal{\textbf{W}}^{r+1}_{i}\leftarrow\{\mathcal{\textbf{W}}^{r+1}_{u,i},\mathcal{\textbf{W}}^{r+1}_{s,i}\}.$ (4)
In Algorithm 2, the resource allocation scheme and cut layer decision based on idle resource states of selected users are shown in Step $5$. In this section, we focus on synchronous federated learning, where the aggregation of all selected user-side ML models and server-side models can only be conducted after one-round training is finished.
Figure 3: This figure illustrates the federated aggregation, where the previous global ML model is $\mathcal{\textbf{W}}^{r}$, the aggregated global ML model is $\mathcal{\textbf{W}}^{r}_{*}$ and the updated global ML model is $\mathcal{\textbf{W}}^{r+1}$
Input: The number of model aggregation round $R$, global learning rate $\eta$, local learning rate $\rho$, and user $i$’s number of local epochs $\epsilon_{i}$;
Output : Final Updated global model $\textbf{W}^{R+1}$;
1 Initialize the global model parameters W;
2 for $r=1,2,...,R$ do
3      The server randomly selects users $S$ to participate the $r$-th round training;
4      The server acquires states of available computing resource $c^{*}$, communication resource (uplink data rate) $b^{*}$, memory $m^{*}$, and storage space $s^{*}$ from $S$;
5      The server allocates its resources $C^{*}$ and $B^{*}$, and determine cut layer $l^{*}$ based on $c^{*}$ and $b^{*}$;
6      for selected user $i=1,2,...,|S|$ do
7           Send $\mathcal{\textbf{W}}_{u,i}^{r}$ to user $i$ based on $l_{i}$ and $\mathcal{\textbf{W}}^{r-1}$;
8           for local epoch $e=1,2,...,\epsilon_{i}$ do
[…]
10
11                end for
12               User $i$ sends back the updated $\mathcal{\textbf{W}}_{u,i}^{r}$ to the server;
13
14                end for
15               Server-side model update: $\mathcal{\textbf{W}}_{s}^{r+1}=\mathcal{\textbf{W}}_{s}^{r}-\eta\sum_{i}\frac{n_{i}}{N}\nabla\mathcal{\textbf{W}}_{s,i}^{r}$;
16                User-side model update: $\mathcal{\textbf{W}}_{u}^{r+1}=\mathcal{\textbf{W}}_{u}^{r}-\eta\sum_{i}\frac{n_{i}}{N}\nabla\mathcal{\textbf{W}}_{u,i}^{r}$;
17
18                end for
Algorithm 2 Efficient Split Federated Learning
III-B3 Communication Model
Since we only intend to demonstrate the effectiveness of our proposed ML scheme, we will adopt a simple communication system model for our study. For uplink and downlink transmissions, we assume uploading and downloading transmission bandwidth are equal. Specifically, the orthogonal multiple access (OMA) techniques are adopted where each user can be allocated with one orthogonal spectrum band for the needed data transmissions determined by the server (e.g., the base station). The uploading and downloading data rate for user $i$ is given by:
$b_{i}^{u}$ $=B_{i}\log\left(1+\frac{P^{u}_{i}\gamma_{i}^{u}}{B_{i}N_{0}}\right),$ (5)
$b_{i}^{d}$ $=B_{i}\log\left(1+\frac{P^{d}_{i}\gamma_{i}^{d}}{B_{i}N_{0}}\right),$
where the communication bandwidth allocated to the user $i$ is $B_{i}$, and the total available bandwidth is $B$ where $\sum_{i=1}^{S}B_{i}\leq B$, $\gamma_{i}^{u}$ is the uplink channel gain and $\gamma_{i}^{d}$ is the downlink channel gain for user $i$, $P^{u}_{i}$ and $P^{d}_{i}$ are the uplink and downlink transmission powers, respectively, when the resource block $B_{i}$ is used, where $P^{u}_{i}$, $P^{d}_{i}$, $\gamma_{i}^{u}$, and $\gamma_{i}^{d}$ are predetermined constant values. $N_{0}$ is the noise power density. We assume that the transmission environment is stationary during one training round. For example, EDs can be cameras in smart homes, whose deployment locations remain fixed for a relatively long period.
III-B4 Workload and Resource Allocation
To reduce the training workload on user-side EDs, we split ML training into local training and server ML training. Thus, the next problem is how to appropriately determine user-side communication and computing workload, alongside the strategic allocation of computing resources on the server to different virtual servers. Due to the varying sizes of data collected by different users/EDs and the heterogeneous available resources on EDs, simply assigning the same amount of resources and randomly choosing the cut layers for different users will not help the training time and the performance. To maximize the training efficiency, one should address a joint optimization of workload and resource.
We denote the total training time as $T$ and the $r$-th round training time as $T_{r}$. To cope with the aforementioned joint optimization problem, we formulate the total training time as
$\begin{gathered}T=\sum_{r=1}^{R}T_{r}=\sum_{r=1}^{R}\max_{i}T_{r,i},\end{gathered}$ (6)
where $T_{r,i}$ is the $r$-th round training time for user $i$.
Since the times of individual training rounds are independent, minimizing the total training time is equivalent to minimizing the training time for each training round. Thus, in the subsequent development, we will only focus on one round of training and omit the training round index r for notational simplicity. One-round training time is composed of four parts, namely, model distribution time $T^{D}_{i}$, model upload time $T^{U}_{i}$, model aggregation time $T^{agg}$, and training time $T^{e}_{i}$ for local epoch $e$. Thus, the one-round training time can be expressed as
$\begin{gathered}T_{i}=T^{U}_{i}+T^{D}_{i}+{\sum_{e=1}^{\epsilon_{i}}T^{e}_{i}}+T^{agg},\\ T_{i}^{U}=\frac{M^{l_{i}}_{i}}{b_{i}^{u}},~{}T_{i}^{D}=\frac{M^{l_{i}}_{i}}{b_{i}^{d}},\\ \end{gathered}$ (7)
where $b_{i}^{U}$ and $b_{i}^{D}$ are the uplink and downlink data rate for user $i$ as defined in (5). We denote the index of the cut layer for user $i$ as $l_{i}$ and $\epsilon_{i}$ as the number of local epochs for user $i$. The uploading and downloading workload are assumed to be $M^{l}_{i}$, equal to the user-side model size of user $i$. Since each selected user applies the same local dataset to train ML model and the idle resource is also stationary, to simplify the following optimization problem, in this study, we assume that the training times of all epochs are constant in each round of training time.
Since both user-side EDs and the server participate in each epoch model updating, one-epoch training time contains four parts, namely, local/user-side computing time $t_{i}^{c}$, uploading time for activation $t_{i}^{b}$, remote/server-side computing time $t_{i}^{C}$, and downloading time for updated activation $t_{i}^{B}$. Thus, one epoch time can be denoted as
$\begin{gathered}T^{e}_{i}=t^{c}_{i}+t^{b}_{i}+t^{C}_{i}+t^{B}_{i}\\ =\frac{w^{c}_{i}}{c_{i}}+\frac{w^{b}_{i}}{b_{i}}+\frac{W^{c}_{i}}{C_{i}}+\frac{w^{b}_{i}}{B_{i}}\\ =\frac{\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}}{c_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}\cdot n_{i}}}{b^{u}_{i}}+\\ \frac{(D-\sum_{l}{D^{l}_{c}x^{l}_{i}})\cdot n_{i}}{C_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}{b^{d}_{i}}.\\ \end{gathered}$ (8)
The cut layer for user $i$ is $l_{i}\in[1,2,...,L]$, where the whole ML neural network is composed of $L$ layers, and the server-side model and the user-side model for user $i$ are split at the $l_{i}$-th layer. To simplify the notation of workloads, we denote $x^{l}_{i}$ as a $[0,1]$ vector, where $\sum_{l}x^{l}_{i}=1$. The total computing workload for training one sample is $D$ and the user-side computing workload for user $i$ is $w^{c}_{i}=\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}$, where $D^{l}$ is the computing workload for one training sample for user-side EDs when the cutting layer is $l$. The upload and download data size for user $i$ is $w_{i}^{b}=\sum_{l}{{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}$, where $D^{l}_{b}$ is the size of activation data for one training sample while the cut layer is $l$. For user $i$, computing capability is $c_{i}$, uplink transmission rate is $b^{u}_{i}$, and downlink transmission rate is $b^{d}_{i}$. The computing resource that the server allocates to user $i$ is $C_{i}$.
IV Optimization and Solution Approach
The ultimate objective of the ESFL training is to minimize the total training time in each round given by (as we mentioned earlier, we will omit the dependence of the round number $r$ for notational simplicity)
$\begin{gathered}\min T=\min\max_{i}T_{i}.\end{gathered}$ (9)
To run Algorithm 2, we need to consider a few optimization problems for resource allocation, which should be solved during the ESFL training. The joint resource allocation and model splitting in our ESFL is a min-max optimization problem. To linearize the formulated optimization problem, we introduce an auxiliary variable $T_{max}$, which is no less than the training time for the straggler (i.e., the client who takes the longest time to complete one-round training). Thus, our problem can be formulated as
$\begin{gathered}\min_{{x_{i}^{l}},C_{i}}T_{max}\\ s.t.~{}T_{i}\leq T_{max},~{}i\in\{1,...,S\}\\ \sum_{i=1}^{S}C_{i}\leq C_{total},\\ \sum_{l=1}^{L}{x_{i}^{l}}M^{l}_{i}\leq s_{i},\\ \sum_{l=1}^{L}{x_{i}^{l}}m^{l}_{i}\leq m_{i},\\ ~{}\sum_{l=1}^{L}x^{l}_{i}=1,~{}x^{l}_{i}\in\{0,1\}.\\ \end{gathered}$ (10)
The total computing resource owned by the server is $C_{total}$, $M^{l}_{i}$ is the data size of user $i$’s user-side ML model, and $s_{i}$ is the available storage space at user $i$. To compute the user-side model, $m^{l}_{i}$ is the required memory space, and $m_{i}$ is the available memory space for user $i$. We denote $x_{i}=\{x_{i}^{1},...,x_{i}^{L}\}$ as the cut layer indicator vector, where $x^{l}_{i}$ indicates whether the ML model is split at layer $l$, in the sense that $x^{l}_{i}=1$ when $l=l_{i}$, and $x^{l}_{i}=0$ otherwise.
Then, our ESFL optimization problem can be transformed into a mixed-integer non-linear program (MINLP), which is typically NP-hard. To solve the problem efficiently, we decompose it into two subproblems and solve them iteratively. We construct the first subproblem for cut layer decision by treating computing resource allocation as variables:
$\begin{gathered}\min_{x_{i}^{l}}T^{e}_{i}=\frac{\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}}{c_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}\cdot n_{i}}}{b^{U}_{i}}+\\ \frac{(D-\sum_{l}{D^{l}_{c}x^{l}_{i}})\cdot n_{i}}{C_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}{b^{D}_{i}}\\ s.t.~{}\sum_{l=1}^{L}x^{l}_{i}=1,~{}x^{l}_{i}\in\{0,1\},\\ M^{l}_{i}\leq s_{i},\\ m^{l}_{i}\leq m_{i}.\\ \end{gathered}$ (11)
Leveraging the iterative optimization approach, the correlation between different users can be eliminated in the sense that we can focus on solving the first subproblem for each user independently, as shown in (11). This is because the cut layer decision for each user is independent of others when computing resource allocation is given. The resulting subproblem for cut layer decision can be easily solved by a linear programming (LP) solver or exhaustive search with the time complexity reduced from $\mathcal{O}(L^{S})$ to $\mathcal{O}(SL)$.
Based on the determined cut layers, we construct the second subproblem for the resource allocation scheme for computing resources for user $i$ as
$\begin{gathered}\min_{C_{i}}T^{e}_{i}=\frac{w_{c}^{l_{i}}\cdot n_{i}}{c_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{u}_{i}}+\\ \frac{(D-w_{c}^{l_{i}})\cdot n_{i}}{C_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{d}_{i}}\\ s.t.~{}\sum_{i=0}^{S}C_{i}\leq C_{total},\end{gathered}$ (12)
where communication and computing workload for all users are constants since the cut layers $l_{i}$ are predetermined by solving the previous subproblem.
The alternative optimization algorithm is shown in Algorithm 3. For the notational convenience, we use $\{C_{*}\}$ to denote $\{C_{1},C_{2},...,C_{S}\}$ and $\{l_{*}\}$ denote $\{l_{1},l_{2},...,l_{S}\}$.
1 Initialization: Allocating identical computing resource to all users, $C_{i}=\frac{C_{total}}{|S|}$;
2 while $C_{i}^{n-1}=C_{i}^{n}$ do
3      Obtain the optimal $\{l_{*}\}$ of subproblem for cut layer decision for given $\{C_{*}\}$ by solving (11);
4      Obtain the optimal $\{C_{*}\}$ of subproblem for resource allocation for the given cut layer decision $\{l_{*}\}$ by solving (12);
5      end while
     Output : $\{l_{*}\}$, $\{C_{*}\}$ for problem (10)
Algorithm 3 Alternative Optimization
V Experiments
In this section, we demonstrate that our ESFL can inherently offer similar performance under the same resource limitation for all users, while significantly reducing total training time (time efficiency). We then show the superior training performance with limited resources and limited training time (model performance). Finally, we validate our iterative optimization approach under various system settings.
V-A Experimental Setup
For all following experiments, we evaluate the performance on image classification tasks over the common dataset CIFAR-10, and leverage VGG13, VGG16 and VGG19 [25] framework as the neural network architecture to implement the distributed applications. We compare our ESFL with the alternatives such as FedAVG (FL), original split learning (SL), splitfed learning (SFL).
Dataset CIFAR-10 [26] contains $50,000$ training samples and $10,000$ testing samples in $10$ classes, with $6,000$ images per class. We assume that there are $100$ users participating in the whole training, while only $10$ users will be randomly selected to join one-round training. Under the assumption that all users’ data samples are independently and identically distributed (IID), the data is shuffled and then partitioned into $100$ clients, each receiving $500$ samples.
Training Configuration We use a distributed machine learning framework, similar to federated learning, which has several learning hyperparameters including local learning rate $\rho_{r}$, where $\rho_{0}=0.01$ and $\rho_{r}$ will be decaying as the round number $r$ increases, and the constant number of local epochs $\epsilon_{i}=5$. Moreover, we introduce a global learning rate $\eta=0.5$ to control the global model updating pace. According to the experimental results, when choosing a mini-batch size of $32$, we can obtain a well trained model.
Neural Network Architecture We deploy the VGG19 network [25] as the training model, which primarily consists of convolutional layers (CONV), fully-connected layers (FC), and softmax layer (SoftMax). We resize the input layer of the original VGG13, VGG16 and VGG19 from $224\times 224$ to $32\times 32$ to fit the CIFAR-10 dataset. The mini-batch size is set to 32. We present VGG19 architecture and workload of each layer in Table I.
TABLE I: VGG19 Network Architecture and Parameters
Layer Layer size FP FLOPs Activation
(MBs) (MBs) (MBs)
CONV1 0.0017 1.796 0.0655
CONV2 0.0369 37.749 0.0328
CONV3 0.0737 18.874 0.0328
CONV4 0.147 37.749 0.0164
CONV5 0.295 18.874 0.0164
CONV6 0.590 37.749 0.0164
CONV7 0.590 37.749 0.0164
CONV8 0.590 37.749 0.0082
CONV9 1.180 18.874 0.0082
CONV10 2.359 37.749 0.0082
CONV11 2.359 37.749 0.0082
CONV12 2.359 37.749 0.0020
CONV13 2.359 9.437 0.0020
CONV14 2.359 9.437 0.0020
CONV15 2.359 9.437 0.0020
CONV16 2.359 9.437 0.0010
FC1 102.760 2.097 4.08E-5
FC2 16.777 0.524 4.08E-5
FC3 4.096 0.131 1E-5
[…]
V-B Model Performance
The key hyperparameter in our ESFL that affect the final convergence performance such as testing accuracy and loss is the number of training rounds, since we leverage FedAVG for all distributed ML algorithms except SL. For a fair comparison, we set the training threshold for VGG13 to $88\%$, VGG16 to $87.5\%$ and VGG19 to $86.5\%$ testing accuracy based on the worst converged accuracy. Three distributed ML algorithms (FL, SFL, ESFL) achieve the expected convergence performance at the $1500$-th training round, and SL achieves the expected convergence performance at the $200$-th training round. The testing performance results are shown in Figure. 4.
(a) Testing accuracy and loss of VGG13, VGG16 and VGG19 using FL
(b) Testing accuracy and loss of VGG13, VGG16 and VGG19 using SL
Figure 4: Testing accuracy and loss over CIFAR-10 testing dataset for FL, SFL, ESFL and SL using three different NN (VGG13, VGG16 and VGG19). Fair comparison are guaranteed by the required training rounds to achieve the convergence threshold.
The rationals for choosing IID configuration rather than Non-IID in our model training process is to mitigate the impact of data distribution heterogeneities for fair training performance comparisons across different distributed ML frameworks.
V-C Time Efficiency
In our simulation, at each round, the server randomly selects $10\%$ users (the selected users) from available users to join one-round training. Since we assume that the server possesses sufficient but limited computing resources, in this experiment, the training server is installed with an A100 GPU with $130$ teraFLOPs (TFLOPs) computing capability and $128$ GigaBytes (GBs) memory space. We compare the time efficiency of our ESFL with original federated learning (FL) [10], split learning [27] and splitfed learning (SFL) [11].
V-C1 Resource Limitation
We separate the impacts of user-side communication and computing resource limitation by simulating four resource settings shown in Table II: Both Poor (BP) indicates that both communication and computing resources are highly limited at EDs, Poorcom Richcmp (PR) indicates that communication resources are highly limited while computing resources are slightly limited (five times larger than that for the highly limited case), Richcom Poorcmp (RP) indicates that communication resources are slightly limited while computing resources are highly limited, and Both Rich (BR) indicates that both communication and computing resources are slightly limited. The selection process for each user-side resource setting is similar to Table II. For learning algorithmic implementation, we use original FL and SL, and SFL, which is similar to that for ESFL introduced in Section III.
TABLE II: Communication and computing resource settings
Communication(KBps) Computing(TFLOPs)
BP [$10$, $15$, $20$, $25$] [$1.3$, $1.95$, $2.6$, $3.25$]
PR [$10$, $15$, $20$, $25$] [$6.5$, $9.75$, $13$, $16.25$]
RP [$50$, $75$, $100$, $125$] [$1.3$, $1.95$, $2.6$, $3.25$]
BR [$50$, $75$, $100$, $125$] [$6.5$, $9.75$, $13$, $16.25$]
Table III presents the average one-round training and communication time and one-round communication time, for different NNs under different resource scenarios using FL, SL, SFL and efficient split federated learning ESFL, respectively. Figure. 5 shows the allocation results of user-side training workload represented as cutting-layer distributions. From cutting-layer distributions, as the user-side resource becomes more sufficient, our ESFL applies more identical cutting-layer distributions strategies for all NNs. For the results in those two table, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances. These advantages often stem from the dynamics between user-side computing and communication resource. In scenarios where local computing resource are poor whereas communication resource are rich (RP), ESFL remains less layers of user-side models leveraging more on server-side computing power. Conversely, under the PR scenario, ESFL mitigates these limitations by remaining more layers of user-side model to rely less on communication. Comparing one-round training and communication time of SFL and FL, an intriguing phenomenon emerges. Although SFL leverages the server-side resource to accelerate training, improper user-side workload allocation (model separation) and server-side resource allocation lead to decreased time efficiency.
TABLE III: One-round training and communication time for different resource settings
NNs Training and communication time(s) Communication time(s)
FL SL SFL ESFL FL SL SFL ESFL
BP VGG13 $40.916$ $226.444$ $42.501$ $28.583$ $11.182$ $123.365$ $21.968$ $8.726$
VGG16 $52.960$ $254.787$ $48.476$ $31.125$ $11.599$ $115.900$ $20.284$ $12.828$
VGG19 $63.096$ $282.426$ $53.231$ $31.128$ $10.200$ $107.575$ $18.253$ $16.216$
PR VGG13 $18.351$ $148.283$ $27.775$ $8.245$ $13.300$ $127.126$ $23.043$ $2.461$
VGG16 $20.933$ $143.427$ $27.065$ $10.242$ $13.401$ $114.513$ $20.616$ $2.858$
VGG19 $23.700$ $141.961$ $27.532$ $12.901$ $14.017$ $105.299$ $18.927$ $3.433$
RP VGG13 $34.128$ $130.896$ $29.229$ $19.297$ $4.238$ $50.829$ $8.430$ $10.961$
VGG16 $45.713$ $156.239$ $35.828$ $19.657$ $4.292$ $45.390$ $7.585$ $11.135$
VGG19 $58.275$ $181.994$ $42.161$ $19.855$ $4.566$ $42.756$ $7.019$ $11.123$
BR VGG13 $10.937$ $73.141$ $14.334$ $6.961$ $5.005$ $51.208$ $9.202$ $0.989$
VGG16 $12.929$ $75.028$ $15.194$ $8.671$ $4.674$ $45.801$ $8.118$ $2.621$
VGG19 $15.968$ $78.177$ $16.220$ $9.470$ $5.428$ $42.736$ $7.549$ $3.184$
(a) Layers distributions of VGG13
(b) Layers distributions of VGG16
(c) Layers distributions of VGG19
Figure 5: Cutting-layer distributions (user-side workloads allocation) of three NNs under four different resource limitations using ESFL algorithm.
Table IV indicates the total training and communication time for four ML algorithms to achieve the expected convergence performance. In the context of contrasting FL and SFL, it is imperative to acknowledge that although SFL, similar to ESFL, utilizes server-side computing resource, its overall performance is significantly influenced by by the harmonization of user-side workload and server-side resource allocation. The lack of effective resource allocation strategies can result in inferior performance in SFL when compared to FL. Especially in RP and BR, when communication resource are notably limited in EDs, FL outperforms SFL. This performance discrepancy can be attributed to the lack of effective allocating strategies, which results in SFL demonstrating inferior performance when compared to FL. Nevertheless, our ESFL exhibits a significant increase in efficiency compared to both FL and SFL across all tested scenarios. This provides the evidence that a well-conceived strategy for workload and resource allocation can markedly enhance the efficiency of the whole training process. While the model used here and other ML algorithms are not the state-of-the-art for this task, it does provide sufficient evidence to show that our ESFL can significantly reduce training latency and improve training efficiency by considering user-side resource heterogeneity.
TABLE IV: Total training and communication time for different resource settings
NNs Training and communication time(s) Communication time(s)
FL SL SFL ESFL FL SL SFL ESFL
BP VGG13 $61,374$ $45,288$ $63,751$ $42,874$ $16,773$ $24,673$ $32,952$ $13,089$
VGG16 $79,440$ $50,957$ $72,714$ $46,687$ $17,398$ $23,180$ $30,426$ $19,242$
VGG19 $94,644$ $56,485$ $79,725$ $46,692$ $15,300$ $21,515$ $27,379$ $24,324$
PR VGG13 $27,526$ $29,656$ $41,662$ $12,367$ $19,950$ $25,425$ $34,564$ $3,691$
VGG16 $31,399$ $28,685$ $41,298$ $15,363$ $20,101$ $22,902$ $30,924$ $4,287$
VGG19 $35,550$ $28,392$ $41,298$ $19,351$ $21,025$ $21,059$ $28,390$ $5,149$
RP VGG13 $51,192$ $26,179$ $43,843$ $28,945$ $6,357$ $10,165$ $12,645$ $16,441$
VGG16 $68,569$ $31,247$ $53,742$ $29,485$ $6,438$ $9,078$ $11,377$ $16,702$
VGG19 $87,412$ $36,398$ $63,241$ $29,782$ $6,849$ $64,134$ $10,528$ $16,684$
BR VGG13 $16,405$ $14,628$ $21,501$ $10,441$ $7,507$ $10,241$ $13,803$ $1,483$
VGG16 $19,393$ $15,005$ $22,791$ $13,006$ $7,011$ $9,160$ $12,177$ $3,931$
VGG19 $23,952$ $15,635$ $24,330$ $14,205$ $8,142$ $8,547$ $11,323$ $4,776$
V-C2 Resource Heterogeneity
In this section, we assume there exist four different communication and computing resource settings to evaluate the efficiency of our ESFL algorithm under different heterogeneous scenarios. The detail of our resource simulation setting is shown in Table V, where in every training round, the available communication and computing conditions of the selected users are randomly chosen from resource options in one scenario. In particular, considering the fairness, the average resource amounts in different scenarios are equal, and only resource distributions are dissimilar to simulate different resource heterogeneity. Four heterogeneous scenarios are considered: Small Heterogeneity (SH), implying that both communication and computing resource heterogeneity at EDs is small, Smallcom Largecmp (SL), implying that communication heterogeneity is small while computing resource is large, Largecom Smallcmp (LS), implying that communication heterogeneity is large while computing resource is small, and Large Heterogeneity (LH), implying that both communication and computing resource heterogeneities are large. For instance, in SH, each user will randomly choose one communication condition from [$10$, $15$, $20$, $25$] kiloBytes (KBps) and one computing condition from [$1.3$, $1.95$, $2.6$, $3.25$] TFLOPs as their available resources, and the server will base on the resource information of the selected users to allocate appropriate server-side computing resource and make user-side cutting layer decision for all selected users. To simulate the training workload heterogeneity, we assume all available users have heterogeneous but constant amounts of data samples, which are chosen from [$200$, $400$, $600$, $800$].
TABLE V: Communication and computing resource settings
Communication(KBps) Computing(TFLOPs)
SH [$10$, $15$, $20$, $25$] [$1.3$, $1.95$, $2.6$, $3.25$]
SL [$10$, $15$, $20$, $25$] [$0.65$, $1.3$, $2.6$, $4.55$]
LS [$5$, $10$, $20$, $35$] [$1.3$, $1.95$, $2.6$, $3.25$]
LH [$5$, $10$, $20$, $35$] [$0.65$, $1.3$, $2.6$, $4.55$]
Table VI presents the average one-round training and communication time and one-round communication time, for different NNs under different resource scenarios using FL, SL, SFL and ESFL, respectively. For the results in those two table, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances and is least affected by resource heterogeneity, where both communication and computing heterogeneity seriously impact the training efficiencies of the other three ML algorithms. It is noteworthy that as the distribution of resources approaches a state of greater uniformity (SH), the gap of time efficiency between SFL and ESFL is decreased. Conversely, with an increase in resource heterogeneity (LH), the performance differential between ESFL and SFL widens notably, which highlight the robustness of ESFL in diverse resource environments. For instance, the training latency of SFL in LH is increased by nearly two times compared with that in SH while training latencies of ESFL are nearly the same in all scenarios. The user-side training workload allocation results (cutting-layer distributions) is shown in Figure. 6. ESFL algorithm demonstrates a trend implementing more uniform cutting-layer distributions across all NNs, under more identical resource distributions.
TABLE VI: One-round training and communication time for different heterogeneities
NNs Training and communication time(s) Communication time(s)
FL SL SFL ESFL FL SL SFL ESFL
SH VGG13 $50.706$ $362.412$ $66.277$ $34.043$ $22.853$ $257.278$ $45.766$ $6.170$
VGG16 $64.136$ $362.872$ $67.640$ $40.658$ $24.544$ $225.720$ $40.473$ $11.093$
VGG19 $76.288$ $381.228$ $69.879$ $43.074$ $24.063$ $210.248$ $35.558$ $22.941$
SL VGG13 $78.595$ $418.746$ $83.431$ $55.912$ $19.913$ $251.753$ $44.067$ $17.867$
VGG16 $103.394$ $457.030$ $92.968$ $58.576$ $21.957$ $229.795$ $39.047$ $26.674$
VGG19 $126.801$ $502.072$ $105.241$ $61.022$ $22.728$ $214.884$ $36.744$ $31.815$
LS VGG13 $79.478$ $518.936$ $109.171$ $37.625$ $53.200$ $415.204$ $90.510$ $9.832$
VGG16 $87.448$ $512.143$ $109.142$ $48.904$ $51.658$ $369.021$ $82.441$ $10.685$
VGG19 $95.622$ $522.882$ $106.500$ $54.211$ $49.811$ $347.786$ $74.140$ $19.720$
LH VGG13 $91.815$ $565.541$ $123.469$ $61.865$ $41.606$ $399.900$ $86.496$ $14.211$
VGG16 $117.439$ $608.637$ $130.643$ $73.708$ $40.871$ $386.282$ $80.686$ $17.929$
VGG19 $145.186$ $601.846$ $130.470$ $75.520$ $44.473$ $323.242$ $66.901$ $42.704$
(a) Layers distributions of VGG13
(b) Layers distributions of VGG16
(c) Layers distributions of VGG19
Figure 6: Cutting-layer distributions (user-side workloads allocation) of three NNs under four different resource heterogeneities using ESFL algorithm.
Table VII indicates the total training and communication time under different RH. It can be seen from the results that our ESFL method is significantly more efficient compared with the original FL, SL and SFL in most scenarios, only except VGG13 in SL. Comparing communication time of FL in SL and LS, the performance of FL is markedly impacted by the RH. However, ESFL capitalizes on these heterogeneities through strategic allocation: in the environment with low communication heterogeneity (SL), it increases the user-side communication workload while decreases computing workload; in the environment with high communication heterogeneity (LS), it conversely allocates more computing workload to EDs. Therefore, our approach optimizes the time efficiency across diverse scenarios by adaptively leveraging RH.
TABLE VII: Total training time to achieve the convergence performance for different heterogeneities
NNs Training and communication time(s) Communication time(s)
FL SL SFL ESFL FL SL SFL ESFL
SH VGG13 $76,059$ $72,482$ $99,415$ $51,064$ $34,279$ $51,455$ $68,649$ $9,255$
VGG16 $96,204$ $72,574$ $101,460$ $60,987$ $36,816$ $45,144$ $60,709$ $16,639$
VGG19 $114,432$ $72,574$ $104,818$ $64,611$ $36,816$ $42,049$ $53,337$ $34,411$
SL VGG13 $117,892$ $83,749$ $125,146$ $83,868$ $29,869$ $50,350$ $66,100$ $26,800$
VGG16 $155,091$ $91,406$ $139,452$ $84,217$ $32,935$ $45,959$ $58,570$ $40,011$
VGG19 $190,201$ $100,414$ $157,861$ $91,533$ $34,092$ $42,976$ $55,122$ $47,722$
LS VGG13 $119,217$ $103,787$ $163,756$ $56,437$ $79,800$ $83,040$ $135,765$ $14,748$
VGG16 $131,232$ $102,428$ $163,713$ $73,356$ $77,487$ $73,804$ $123,661$ $16,027$
VGG19 $143,433$ $104,576$ $159,750$ $81,316$ $74,716$ $69,557$ $111,210$ $29,580$
LH VGG13 $137,722$ $113,108$ $185,203$ $92,797$ $62,409$ $79,980$ $129,744$ $21,316$
VGG16 $176,158$ $121,727$ $195,964$ $110,562$ $61,306$ $77,256$ $121,029$ $26,893$
VGG19 $217,779$ $120,369$ $195,705$ $113,280$ $66,709$ $64,648$ $100,351$ $64,056$
V-D Resource Allocation Convergence Analysis
From Section IV, the joint resource allocation and model splitting problem has been decomposed into two subproblems due to the time complexity, and we leverage an alternative optimization approach to solve those subproblems. To evaluate the convergence of our alternative method, we simulate iterative results at four users’ scales, 100 users, 200 users, 400 users and 800 users. The simulation result in Figure. 7 shows that even for the largest user scale (800 users), our alternative approach only needs $10$ iterations to achieve convergence, while when the user scales are small, it only requires a few iterations to achieve convergence.
Figure 7: Average one-round training time using iterative optimization.
VI Conclusion
In this paper, we have designed ESFL, a novel distributed training approach that tackles the resource heterogeneity inherent in both federated learning and split learning. Unlike previous methods in addressing data heterogeneity in FL, we have provided a new perspective by allocating appropriate server-side resources and user-side workload to effectively address the straggler problem in the synchronous FL framework. By evaluating the training efficiency for different ML algorithms under different heterogeneous scenarios, we have performed extensive analysis and demonstrated the superiority of our proposed ESFL.